OpenStudy (anonymous):

Simplify 4ab * √(2b) - 3a * √(18b³) + 7ab * √(6b)

OpenStudy (abb0t):

You can simplify the \(\sqrt{18b^3}\). Notice that you can break it down to \(\sqrt{6 \times 3 (b \times b \times b)} = \sqrt{(3 \times 3 \times 2)(b \times b \times b)}\). And remember also when you have pairs, you can take one ONE of the pairs. For instance: \(\sqrt{a^3} = \sqrt{a \times a \times a} = a \sqrt{a}\). You square roots took out a pair of a and left one a inside because you can only take out by pairs.

OpenStudy (anonymous):

Thanks i got it